OpenStudy (anonymous):

A 0.684 g sample of carbon was found to contain 3.81 x 1020 atoms of 13C. What is the natural abundance of this isotope? Select one: a. 11.1 % b. 1.01 % c. 0.111 % d. 1.11 %

OpenStudy (unklerhaukus):

total number of atoms=\[\Big\{0.684[\text g]\text C\Big\}\div\left\{12\frac{[\text g]\text C}{[\text {mol}]}\right\}\times\Big\{6.022\times10^{23}\frac{[\text{molecules}]}{[\text{mol}]}\Big\}\]

OpenStudy (unklerhaukus):

The total number of \(\text C\) atoms \[\Big\{0.684[\text g]\text C\Big\}\div\left\{12.01\frac{[\text g]\text C}{[\text {mol}]}\right\}\times\Big\{6.022\times10^{23}\frac{[\text{atoms}]}{[\text{mol}]}\Big\}\] Divide the number of \(^{13}\text C\) by the total number of \(\text C\) atoms and multiply by 100% to find the percentage abundance